OpenStudy (clarence):

If A is an invertible matrix then the equation Ax = b may have two distinct non zero solutions for a non zero vector b. True or False?

OpenStudy (clarence):

I personally think that it's false, but confirmation is always nice to have.

ganeshie8 (ganeshie8):

why do you think it is false

OpenStudy (irishboy123):

.

OpenStudy (clarence):

Well...

OpenStudy (clarence):

If there exist two different solutions, for example Ax1 = Ax2 = b, then A(x1−x2) = 0 with x1 − x2 ≠ 0. Wouldn't that follow up to be saying that there are infinite solutions? If A were invertible, you could write 0 = A^(−1) (0) = A^(−1) (A(x1−x2)) = x1−x2, but that's impossible because x1≠x2, hence A cannot be invertible?

OpenStudy (clarence):

If that made any sense ~

ganeshie8 (ganeshie8):

That looks a lot better than what I had : \(A\) is invertible, so \(Ax=b\implies x=A^{-1}b\) since the inverse of a matrix is unique (when it exists), the solution is unique.

OpenStudy (clarence):

Thank you!!

ganeshie8 (ganeshie8):

np